# thanks: These authors contributed equally to this workthanks: These authors contributed equally to this work

### v1

@@ Abstract
The Wineland parameter aims at detecting metrologically useful entangled states, called spin-squeezed states, from expectations and variances of total angular momenta. Efficient strategies for estimating this parameter in practice have yet to be determined. We here formulate the question of whether a state is spin-squeezed according to the Wineland parameter as an hypothesis test and assess the relevance of experimental outcomes by deriving upper and lower bounds on the p-value. By applying our statistical test to data obtained in multiple experiments, we are unable to reject the hypothesis that non-spin squeezed states were measured with a p-value of 5% or less in most cases. We also find an explicit non-spin squeezed state according to the Wineland parameter reproducing most of the observed results with a p-value exceeding 5%. Our results provide a rigorous method to establish robust statistical evidence of spin squeezing from the Wineland parameter in future experiments, accounting for finite-size effects.
@@
$\xi^{2}_{W}=\frac{N\text{ Var}(J_{\vec{n}_{\perp}})}{|\langle J_{\vec{n}}\rangle|^{2}}<1$ (2)
@@
is said to be (metrologically) spin-squeezed along the $\perp$ direction. Here, $\langle J_{\vec{n}}\rangle$ denotes the expectation value of the spin operator along the direction $\vec{n}$ and Var$(J_{\vec{n}_{\perp}})$ is the variance of the spin in a direction $\vec{n}_{\perp}$ orthogonal to $\vec{n}$. $\xi^{2}_{W}$ is called the Wineland parameter. Put simply, $\xi^{2}_{W}<1$ witnesses a class of quantum states where the spin variance is reduced in one direction while being increased in the orthogonal direction [23, 24]. In a Ramsey experiment, $\xi^{2}_{W}$ is established as the ratio of phase sensitivity of the state of interest $\hat{\rho}$ compared to a coherent spin state. Hence $\xi^{2}_{W}<1$ witnesses states having the potential to beat the standard quantum limit. The advantage of spin-squeezed states in metrology is due to quantum correlations between spins [24]. Specifically $\xi^{2}_{W}<1$ witnesses entanglement [25]. The ability to observe multiple spin states that are both metrologically useful and entangled through collective spin measurements has driven numerous experimental implementations with various physical plateforms involving up to almost a million spins [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22]. However, the exact values of the variance and expectation of total angular momenta involved in the Wineland parameter cannot be obtained from a finite number of measurement repetitions. Hence, concluding that spin-squeezed states have been created and detected in such systems requires careful consideration due to the limitations of finite statistics. This challenge was highlighted in Refs. [26, 27], which aimed to draw conclusions about Bell correlations in a many-spin system. In this letter, we present a rigorous method to establish robust statistical evidence of spin squeezing from the Wineland parameter, accounting for finite-size effects.
@@
To set up the statistical test, the hypotheses need to be formulated. As a first step, we focus on the following linearized form of the Wineland parameter
$\Gamma=N(\Delta Q_{{\vec{n}_{\perp}}})^{2}-\langle Q_{\vec{n}}\rangle^{2}$ (3)
where $Q_{i}=2J_{i}/N$ is the normalized angular momemtum operator in a direction specified by $i$. In the asymptotic limit (infinite number of measurement repetitions), a spin system is considered spin-squeezed if the corresponding state satisfies $\Gamma<0$. Therefore, we define the null hypothesis $H_{0}$ as “the state of interest satisfies $\Gamma\geq 0$”. At first glance, the sample variance and sample mean are natural choices as estimators of $(\Delta Q_{i})^{2}$ and $\langle Q_{i}\rangle^{2}$. However, this results in an estimator for $\Gamma$ that is a nonlinear function of independent random variables, creating challenges in accurately estimating the number of experimental repetitions required to reject the null hypothesis $H_{0}$ without overestimation, see Supplemental Material Sec. I and IV. In this work, we instead introduce the following family of criteria
$\Gamma_{c}=N\langle Q^{2}_{\vec{n}_{\perp}}\rangle-f_{\alpha,\beta}(\langle Q_{\vec{n}_{\perp}}\rangle,\langle Q_{\vec{n}}\rangle).$ (4)
Here, $f_{\alpha,\beta}(x,y)=2\alpha Nx+2\beta y-N\alpha^{2}-\beta^{2}$ is the tangent plane of $h(x,y)=Nx^{2}+y^{2}$ at the point $c=(\alpha,\beta)\in[-1,1]$ and forms an affine lower bound for $h(x,y)$. Therefore, for any given point $(\alpha,\beta)$, $\Gamma_{c}$ serves as a linear upper bound of $\Gamma$, that is, $\Gamma\leq\Gamma_{c}$, and a quantum state satisfying $\Gamma_{c}<0$ is classified as spin-squeezed.
@@
Figure 1: Experimental achievements together with their statistical analysis. The experiments that have been reported in the litterature are identified by a square whose position specifies the total number of experimental repetitions ($M$) that has been realized and the system size ($N$). The color of the square gives the observed value of the Wineland parameter. Every square is accompanied by a number used to identify the reference as specified below the graph. The significance of the experimental outcomes are quantified by the number of measurements that are necessary and sufficient to reach a p-value of 5%. Namely, the upper triangle (with the vertex pointing down) gives the total number of measurement repetitions that is sufficient to reject the hypothesis that non-spin squeezed states were measured. The lower triangle (with the vertex pointing up) specifies the number of measurements that is required in order to ensure that a specific non-spin squeeezed state (according to the Wineland parameter) cannot reproduce the observed statistics.
Let $\widetilde{\Gamma}_{{c}}$ be an estimator for $\Gamma_{c}$ and $\gamma_{c}<0$ the observed value in an actual experiment. The significance of the experimental outcome is given by the p-value $P(\widetilde{\Gamma}_{c}\leq\gamma_{c}|H_{0})$. Specifically, If $P(\widetilde{\Gamma}_{c}\leq\gamma_{c}|H_{0})$ is less than a predetermined significance level, such as 0.05, the null hypothesis $H_{0}$ can be rejected and we conclude that the system of interest is spin squeezed.
@@
For the estimator, we consider the following sum of identical and independent random variables
$\widetilde{\Gamma}_{c}=\sum_{i=1}^{M/2}\Gamma_{c,i}$ (5)
with
$\Gamma_{c,i}=N\left(q^{(i)}_{\vec{n}_{\perp}}\right)^{2}-f_{\alpha,\beta}\left(q^{(i)}_{\vec{n}_{\perp}},q^{(i)}_{\vec{n}}\right).$ (6)
Here, $q^{(i)}_{\vec{n}_{\perp}}$ $(q^{(i)}_{\vec{n}})$ is the outcome of the measurement operator $2J_{\vec{n}}^{(i)}/N$, $(2J_{\vec{n}_{{\perp}}}^{(i)}/N)$ for experimental round $i$ and $M/2$ is the number of experimental repetitions in each measurement directions ($M$ is the total number of experimental repetitions). To determine the significance level without presuming the probability distribution of the estimator, we employ the Bernstein inequality to bound $P(\widetilde{\Gamma}_{c}\leq\gamma_{c}|H_{0})$, which gives (see Supplemental Material, Sec. II)
$P(\widetilde{\Gamma}_{c}\leq\gamma_{c}|H_{0})\leq\exp{\dfrac{\gamma^{2}_{c}M/2}{2\Gamma^{c}_{1}\Gamma^{c}_{0}+2(\Gamma^{c}_{1}-\Gamma^{c}_{0})\gamma_{c}/3}}.$ ) . (7)
$\Gamma^{c}_{1}$ and $\Gamma^{c}_{0}$ are the maximal and minimal values of $\Gamma_{c,i}$, which are given by
$\Gamma^{c}_{1}$ $=N(1+|\alpha|)^{2}+(1+|\beta|)^{2}-1,$ (8)
$\Gamma^{c}_{0}$ $=(1-|\beta|)^{2}-1.$
An upper bound on the p-value can thus be obtained in practice by first extracting the value $\gamma_{c}$ that the estimator $\tilde{\Gamma}_{c}$ takes on the observed statistics and then deducing the right hand side of Eq. (7) from the number of experimental repetitions $M$ by fixing $\alpha$ and $\beta$. Ultimately, this upper bound can be tightened by minimizing the right hand side of Eq. (7) over $\alpha$ and $\beta$.
@@
To estimate what could have been achieved in previously implemented experimental realizations, we extracted the value of $\gamma_{c}$ from Refs. [11, 15, 16, 17, 14, 18, 19, 20, 21, 22, 12, 13] by assuming that $\sum_{i=1}^{M/2}q^{(i)}_{\vec{n}_{\perp}}=0$. We increase $M$, optimizing $(\alpha,\beta)$, until the upper bound on the p-value achieves $5$%. We then execute numerical simulations to show that the assumption $\sum_{i=1}^{M/2}q^{(i)}_{\vec{n}_{\perp}}=0$ is conservative, i.e. it minimizes the number of experimental repetitions that are needed to achieve a given upper bound on the $p$-value, see Supplemental Material Sec. III. The value of $M$ that we find is reported in Fig. 1 (triangle with the vertex pointing down) as a function of the system size for each experiment separately. The actual number of experimental repetitions conducted is also indicated (square). Except for the experiment involving 16 spins [^1] In Ref. [13], the authors specify that each measurement point reported in their paper is an average of 50 to 600 repetitions of experimental realizations. The position of the corresponding square in Fig. 1 assumes that 1200 repetitions in total were performed, as two measurements directions are involved . Actually, 728 realizations are enough to achieve an upper bound on the p-value of 5%. reported in Ref. [13] and 470 spins [^2] In Ref. [15], a tomography was performed using a total of 32500 experimental realizations. The position of the corresponding square in Fig. 1 hence assumes that 32500 realizations were performed. Actually, 19970 realizations (9985 in each measurement direction) ensure a significance level of 5%. in Ref. [15], we cannot reject the hypothesis that non-spin squeezed states were measured with a p-value less than or equal to 5%. For most experiments, the discrepancy between the actual number of measurements and the number required to achieve an upper bound on the p-value of 5% is at least a factor of 10, and can reach up to $10^{4}$ in the worst case.
@@
The upper bound on the p-value we derived before suggests that the number of experimental repetitions performed may be insufficient to rule out the possibility that non-spin-squeezed states were actually produced in most existing experimental realisations. To substantiate this claim, we consider the following state
@@
where $\kappa=\xi_{W}^{2}\left(\ket{\chi}\right)-N/\langle Q_{\vec{n}}\rangle_{\ket{\chi}}^{2}$ (see Supplemental Material Sec. V). When performing $M$ measurements on $\hat{\rho}$ with $r\leq r_{\text{max}}$, the probability to only sample the squeezed state |$\ket{\chi}$ occurs with a probability $p=r^{M}\leq r_{\text{max}}^{M}=p^{*}$. Hence, $p^{*}$ establishes a lower bound on the p-value of a statistical test designed to reject the null hypothesis $H_{0}$.
@@
We designed a linear estimator associated to the Wineland parameter and use the Bernstein inequality to quantify the significance of experimental data. This defines a proper recipe for estimating the Wineland parameter in practice. The first step consists in extracting the value $\gamma_{c}$ that the estimator $\tilde{\Gamma}_{c}$ takes on the observed statistics using Eq. (5) and then deducing the upper bound on the p-value from the minimization over $\alpha$ and $\beta$ of the right hand side of Eq. (7) while fixing the number of experimental repetitions $M$. The null hypothesis associated to “the state of interest satisfies $\Gamma\geq 0$” can then be rejected if the p-value is below 0.05 or any other desired value.
Our analysis also provides an estimate of the number of measurements which is necessary (but not sufficient) to reject the null hypothesis. From a specific state, we showed that the p-value is lower bounded by $p^{*}=r_{\max}^{M}$ where $r_{\max}$ is specified from the expected values of $\xi_{W}^{2}$ and $\langle Q_{\vec{n}}\rangle$, see Eq. (10). Note that $r_{\max}\geq(\sqrt{N^{2}+4N}-N)/2$ and hence, a number of measurements $M\geq\ln{0.05}/{\ln{[(\sqrt{N^{2}+4N}-N)/2]}}\sim-N\log(0.05)$ ) is necessary to achieve a p-value smaller than 5%.
Note that the p-value bounds we proposed could likely be improved. For instance, using more refined estimators may yield tighter bounds. Additionally, various concentration inequalities could be explored, and although several have already been tested (see Supplemental Material), other choices might offer even stronger bounds. Similarly, while we have tested multiple states to construct the lower bound on the p-value, we cannot rule out the possibility that better states exist for this purpose. Moreover, similar methods could be developed to quantify finite-size effects on other squeezing or entanglement parameters [28]. We leave these extensions for future work. Nonetheless, our results already offer a solid framework for predicting and accurately assessing the outcomes of future experiments focused on spin squeezing.

### v2

@@ Abstract
The Wineland parameter aims at detecting metrologically useful entangled states, called spin-squeezed states, from expectations and variances of total angular momenta. However, efficient strategies for estimating this parameter in practice have yet to be determined and in particular, the effects of a finite number of measurements remain insufficiently addressed. We formulate the detection of spin squeezing as a hypothesis-testing problem, where the null hypothesis assumes that the experimental data can be explained by non-spin-squeezed states. Within this framework, we derive upper and lower bounds on the p-value to quantify the statistical evidence against the null hypothesis. By applying our statistical test to data obtained in multiple experiments, we are unable to reject the hypothesis that non-spin squeezed states were measured with a p-value of 5% or less in most cases. We also find an explicit non-spin squeezed state according to the Wineland parameter reproducing most of the observed results with a p-value exceeding 5%. More generally, our results provide a rigorous method to establish robust statistical evidence of spin squeezing from the Wineland parameter in future experiments, accounting for finite statistics.
@@
$\xi^{2}_{W}=\frac{N\text{ Var}(J_{\vec{n}_{\perp}})}{\langle J_{\vec{n}}\rangle^{2}}<1$ (2)
@@
is said to be (metrologically) spin-squeezed along the $\vec{n}_{\perp}$ direction. Here, $\langle J_{\vec{n}}\rangle$ denotes the expectation value of the spin operator along the direction $\vec{n}$ and Var$(J_{\vec{n}_{\perp}})$ is the variance of the spin in a direction $\vec{n}_{\perp}$ orthogonal to $\vec{n}$. $\xi^{2}_{W}$ is called the Wineland parameter. Put simply, $\xi^{2}_{W}<1$ witnesses a class of quantum states where the collective spin variance is reduced in one direction while being increased in the orthogonal direction [23, 24]. In a Ramsey experiment, $\xi^{2}_{W}$ is established as the ratio of phase sensitivity of the state of interest $\hat{\rho}$ compared to a coherent spin state. Hence $\xi^{2}_{W}<1$ witnesses states having the potential to beat the standard quantum limit. The advantage of spin-squeezed states in metrology is due to quantum correlations between spins [24]. Specifically $\xi^{2}_{W}<1$ witnesses entanglement [25]. The ability to observe multiple spin states that are both metrologically useful and entangled through collective spin measurements has driven numerous experimental implementations involving up to almost a million spins [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22].
Experimentally confirming spin-squeezing has traditionally relied on the error bar associated to the estimation of the Wineland parameter. However, to demonstrate evidence of spin-squeezing, one should quantify the probability that the observed statistics, obtained from a limited number of experimental repetitions, could result from measurements on a state with no spin-squeezing, i.e. a state satisfying $\xi^{2}_{W}\geq 1$. Error bars only quantify the spread of the measurement outcomes. They do not provide information about the likelihood that the observed data could have been produced by non-spin-squeezed states [26]. A more comprehensive statistical framework is thus needed for testing the Wineland criterion in practice. This challenge was highlighted in Refs. [27], where a similar issue emerged in the context of Bell correlation witnesses in many-spin systems.
In this letter, we develop a rigorous statistical framework for establishing robust evidence of spin squeezing based on the Wineland parameter, explicitly accounting for finite-size effects. Our approach focuses on quantifying the probability (p-value) that the observed experimental data could arise from a non-spin-squeezed state, thereby formulating the problem as a hypothesis test. Unlike the approach in Ref. [28], our analysis does not test a specific non-spin-squeezed state and enables the rejection of the hypothesis that any non-spin squeezed states could have produced the observed statistics with a specified p-value. Our findings indicate that most existing experiments do not sufficiently account for finite-size effects, with only a limited number of cases capable of rejecting the hypothesis test. In many situations, an insufficient number of measurements has been performed which result in a high probability that a non-spin-squeezed state could mimic the experimental outcomes, particularly when the total spin number is large.
Traditional approach —
The verification of spin squeezing in experiments is typically based on the statistical estimation of the Wineland parameter $\xi^{2}_{W}$. Conventional approaches involve measuring the total spin $\langle J_{\vec{n}}\rangle$ and the variance $\text{Var}(J_{\vec{n}_{\perp}})$, followed by an evaluation of the uncertainty in $\xi^{2}_{W}$ from its error bar. A state is conventionally classified as spin-squeezed if the estimated value of $\xi^{2}_{W}$ falls below 1 with a sufficiently small error bar. However, error bars do not constitute a rigorous test for statistical significance [26]. In particular, they fail to account for the probability that statistical fluctuations may cause non-spin-squeezed states to satisfy $\xi^{2}_{W}<1$, leading to potential misclassification. To illustrate this risk, we conduct a Monte Carlo simulation by repeatedly measuring a non-spin-squeezed state of $9$ spins with $\xi_{W}^{2}=1$, as detailed in End Matter. The results show that even with a large number of measurements resulting in a small error bar, statistical fluctuations can still cause a non-spin-squeezed state to be mistakenly classified as spin-squeezed.
@@
We here focus on null-hypothesis significance testing, with the p-value quantifying the probability of obtaining a Wineland parameter at least as extreme as the value actually observed, under the null hypothesis $H_{0}$ that the measured state is non-spin-squeezed. A small p-value thus provides strong evidence against $H_{0}$, suggesting that the observed data is unlikely under the assumption that they result from measurements of a non-spin-squeezed state. To ensure the p-value remains independent of any specific quantum state or statistical distribution, we determine an upper bound by considering all possible non-spin-squeezed states.
To formulate the null-hypothesis significance test, we introduce the normalized spin angular momentum operator $Q_{(\cdot)}=2J_{(\cdot)}/N$, define the following parameter
$\Gamma=N\text{Var}(Q_{{\vec{n}_{\perp}}})-\langle Q_{\vec{n}}\rangle^{2}.$ (3)
and use the sample variance and sample mean as unbiased estimators for $\text{Var}(Q_{\vec{n}_{\perp}})$ and $\langle Q_{\vec{n}}\rangle^{2}$. In the asymptotic limit (infinite measurement repetitions), a spin system is considered spin-squeezed if the corresponding state satisfies $\Gamma<0$. While Eq. (3) provides a natural alternative criterion for spin squeezing, it involves nonlinear terms. Existing concentration inequalities struggle to effectively account for these nonlinearities, resulting in an overly conservative upper bound on the p-value. As a result, the predicted number of measurements required to achieve a small p-value is overestimated (see Supplemental Material, Secs. II and V).
To address these challenges, we propose a family of parameters
$\Gamma_{c}=N\langle Q^{2}_{\vec{n}_{\perp}}\rangle-f_{\alpha,\beta}(\langle Q_{\vec{n}_{\perp}}\rangle,\langle Q_{\vec{n}}\rangle),$ (4)
where each element is defined by a fixed value of $c=(\alpha,\beta)$. The function $f_{\alpha,\beta}(x,y)$ is the tangent plan to $Nx^{2}+y^{2}$ at the point $c=(\alpha,\beta)\in[-1,1]$
$f_{\alpha,\beta}(x,y)=2\alpha Nx+2\beta y-N\alpha^{2}-\beta^{2}.$ (5)
Importantly, $f_{\alpha,\beta}(x,y)$ serves as an affine lower bound for $Nx^{2}+y^{2}$ for any $(x,y)$ and $(\alpha,\beta)$. This ensures that $\Gamma_{c}\geq N\langle Q^{2}_{\vec{n}_{\perp}}\rangle-N\langle Q_{\vec{n}_{\perp}}\rangle^{2}-\langle Q_{\vec{n}}\rangle^{2}=\Gamma$ for any choice of $c=(\alpha,\beta)$. As a result, quantum states satisfying $\Gamma_{c}<0$ can be classified as spin-squeezed.
@@
Figure 1: Experimental achievements together with their statistical analysis. The experiments that have been reported in the litterature are identified by a square whose position specifies the total number of experimental repetitions ($M$) that has been realized and the system size ($N$). The color of the square gives the observed value of the Wineland parameter. Every square is accompanied by a number used to identify the reference as specified below the graph. The significance of the experimental outcomes are quantified by the number of measurements that are necessary and sufficient to reach a p-value of 5%. Namely, the upper triangle (with the vertex pointing down) gives the total number of measurement repetitions that is sufficient to reject the hypothesis that non-spin squeezed states were measured. The lower triangle (with the vertex pointing up) specifies the number of measurements that is required in order to ensure that a specific non-spin squeezed state (according to the Wineland parameter) cannot reproduce the observed statistics.
To estimate $\Gamma_{c}$ in practice, we consider $M/2$ independent and identically distributed (i.i.d.) repeated processes ($M$ is the total number of experimental repetitions). In each processes $i$, the system undergoes two independent measurements. In the first measurement, the quantum state is prepared, and a measurement of $Q_{\vec{n}}$ is performed, yielding the outcome $q^{(i)}_{\vec{n}}\in[-1,1]$. In the second measurement, the quantum state is re-prepared, and a measurement of $Q_{\vec{n}_{\perp}}$ is performed, obtaining the outcome $q^{(i)}_{\vec{n}_{\perp}}\in[-1,1]$. Using these measurement outcomes, we define the estimator of $\Gamma_{c}$ for each experimental processes $i$ as the random variable
$\widetilde{\Gamma}_{c,i}=N\left(q^{(i)}_{\vec{n}_{\perp}}\right)^{2}-f_{\alpha,\beta}\left(q^{(i)}_{\vec{n}_{\perp}},q^{(i)}_{\vec{n}}\right).$ (6)
Given the $M/2$ repetitions, $\Gamma_{c}$ can be estimated by using the following statistical estimator
$\widetilde{\Gamma}_{c}=\frac{2}{M}\sum_{i=1}^{M/2}\widetilde{\Gamma}_{c,i}.$ (7)
Let $\gamma_{c}<0$ be the observed value of $\widetilde{\Gamma}_{c}$ from a given experimental realization. We are interested in the p-value, i.e. the probability that the estimator $\widetilde{\Gamma}_{c}$ takes a value equal to or smaller than $\gamma_{c}$, assuming the measurements are performed on non-spin-squeezed states (the null hypothesis $H_{0}$). We name it explicitly as $P(\widetilde{\Gamma}_{c}\leq\gamma_{c}\,|\,H_{0})$. Importantly, $P(\widetilde{\Gamma}_{c}\leq\gamma_{c}\,|\,H_{0})$ must be estimated by considering all possible non-spin-squeezed states and all possible statistical distributions. If $P(\widetilde{\Gamma}_{c}\leq\gamma_{c}\,|\,H_{0})$ is less than a predetermined significance level, such as 0.05, we conclude that the system of interest exhibits spin squeezing.
@@
Concentration inequalities [29, 30] can be used to upper bound $P(\widetilde{\Gamma}_{c}\leq\gamma_{c}\,|\,H_{0})$. As shown in Supplemental Material, Sec. III, the Bernstein inequality for example gives
$P(\widetilde{\Gamma}_{c}\leq\gamma_{c}|H_{0})\leq\exp{\dfrac{\gamma^{2}_{c}M/2}{2\Gamma^{c}_{1}\Gamma^{c}_{0}+2(\Gamma^{c}_{1}-\Gamma^{c}_{0})\gamma_{c}/3}}$ ) (8)
where
$\Gamma^{c}_{1}$ $=\max_{q^{(i)}_{\vec{n}_{\perp}},q^{(i)}_{\vec{n}}}\widetilde{\Gamma}_{c,i}=N(1+|\alpha|)^{2}+(1+|\beta|)^{2}-1,$
$\Gamma^{c}_{0}$ $=\min_{q^{(i)}_{\vec{n}_{\perp}},q^{(i)}_{\vec{n}}}\widetilde{\Gamma}_{c,i}=(1-|\beta|)^{2}-1.$
The right-hand side of Ineq. (8) is solely a function of the parameter $c=(\alpha,\beta)$ and the experimental observation $\gamma_{c}$. It can be minimized over $c=(\alpha,\beta)$ under the constraint that $\gamma_{c}<0$ in order to get the tightest upper bound on the p-value. This defines $p_{\text{opt}}$ as
$p_{\text{opt}}=\inf_{\alpha,\beta\text{ s.t.}\gamma_{c}<0}\exp{\dfrac{\gamma^{2}_{c}M/2}{2\Gamma^{c}_{1}\Gamma^{c}_{0}+2(\Gamma^{c}_{1}-\Gamma^{c}_{0})\gamma_{c}/3}}.$ ) . (9)
Guidelines for robust statistical evidence of spin squeezing in future experiments —
We now have all the ingredients to propose a concrete recipe that can be used to establish a robust statistical evidence of spin squeezing in future experiments. First, we consider $M/2$ measurements of both $Q_{\vec{n}}$ and $Q_{\vec{n}_{\perp}}$. This yields outcomes $q^{(i)}_{\vec{n}}$ and $q^{(i)}_{\vec{n}_{\perp}}$ for $i\in[1,M/2]$. $\gamma_{c}$ is then deduced from the estimator $\widetilde{\Gamma}_{c}$ given in Eq. (6). An upper bound on the p-value is obtained from the minimization given in Eq. (9) over $\alpha$ and $\beta$ such that $\gamma_{c}<0$. A small $p_{\text{opt}}$ provides strong evidence against $H_{0}$, suggesting that the observed data is unlikely under the assumption that they result from measurements of a non-spin-squeezed state.
@@
To estimate what could have been achieved in previously implemented experimental realizations, we calculated the value of $\gamma_{c}$ from Refs. [11, 15, 16, 17, 14, 18, 19, 20, 21, 22, 12, 13] by assuming that $\sum_{i=1}^{M/2}q^{(i)}_{\vec{n}_{\perp}}=0$. We increase $M$, optimizing $(\alpha,\beta)$, until the upper bound on the p-value achieves $5$%. We then execute numerical simulations to show that the assumption $\sum_{i=1}^{M/2}q^{(i)}_{\vec{n}_{\perp}}=0$ is conservative, i.e. it minimizes the number of experimental repetitions that are needed to achieve a given upper bound on the $p$-value, see Supplemental Material Sec. IV. The value of $M$ that we find is reported in Fig. 1 (triangle with the vertex pointing down) as a function of the system size for each experiment separately. The actual number of experimental repetitions conducted is also indicated (square). Except for the experiment involving 16 spins [^1] In Ref. [13], the authors specify that each measurement point reported in their paper is an average of 50 to 600 repetitions of experimental realizations. The position of the corresponding square in Fig. 1 assumes that 1200 repetitions in total were performed, as two measurements directions are involved . Actually, 728 realizations are enough to achieve an upper bound on the p-value of 5%. reported in Ref. [13] and 470 spins [^2] In Ref. [15], a tomography was performed using a total of 32500 experimental realizations. The position of the corresponding square in Fig. 1 hence assumes that 32500 realizations were performed. Actually, 19970 realizations (9985 in each measurement direction) ensure a significance level of 5%. in Ref. [15], we cannot reject the hypothesis that non-spin squeezed states were measured with a p-value less than or equal to 5%. For most experiments, the discrepancy between the actual number of measurements and the number required to achieve an upper bound on the p-value of 5% is at least a factor of 10, and can reach up to $10^{4}$ at the worst case.
@@
The upper bound on the p-value we derived before suggests that the number of experimental repetitions performed may be insufficient to rule out the possibility that non-spin-squeezed states were actually produced in most existing experimental realisations. To substantiate this claim, we follow the approach of Refs. [31, 28] and in particular, we consider the following state
@@
where $\kappa=\xi_{W}^{2}\left(\ket{\chi}\right)-N/\langle Q_{\vec{n}}\rangle_{\ket{\chi}}^{2}$ (see Supplemental Material Sec. VI). When performing $M$ measurements on $\hat{\rho}$ with $r\leq r_{\text{max}}$, the probability to only sample the squeezed state |$\ket{\chi}$ occurs with a probability $p=r^{M}\leq r_{\text{max}}^{M}=p^{*}$. Hence, $p^{*}$ establishes a lower bound on the p-value of a statistical test designed to reject the hypothesis that the measured state is of the form given in Eq. (10[31, 28].
@@
In this paper, we developed a rigorous analytical framework for detecting spin-squeezed states, with a particular focus on the impact of finite measurement repetitions on the detection results. Specifically, by linearizing $\xi_{W}^{2}$, we proposed a family of criteria for identifying spin-squeezed states. Using a hypothesis-testing approach, we derived upper bounds on the p-value, which can be used to provide a reliable statistical evidence of spin squeezing.
The effects of finite measurement repetitions have long been overlooked in spin-squeezing detection experiments. Our hypothesis-testing framework reveals that most existing experiments fail to pass the statistical test due to an insufficient number of measurements. This issue becomes particularly severe when the spin number is large, as the required number of measurements needed to reliably reject the null hypothesis increases significantly, often exceeding the practical capabilities of current experimental setups. However, this does not necessarily mean that spin squeezing is absent, but rather that the available data does not provide strong enough statistical support for claiming spin squeezing is produced.
Our analysis provides an estimate of the number of measurements which is necessary (but not sufficient) to reject the null hypothesis. From a specific state, we showed that the p-value is lower bounded by $p^{*}=r_{\max}^{M}$ where $r_{\max}$ is specified from the expected values of $\xi_{W}^{2}$ and $\langle Q_{\vec{n}}\rangle$, see Eq. (11). Note that $r_{\max}\geq(\sqrt{N^{2}+4N}-N)/2$ and hence, a number of measurements $M\geq\ln{0.05}/{\ln{[(\sqrt{N^{2}+4N}-N)/2]}}\sim-N\log(0.05)$ ) is necessary to achieve a p-value smaller than 5%.
Note that the p-value bounds we proposed in Eq. (9) could likely be improved. For instance, using more refined estimators may yield tighter bounds. Additionally, various concentration inequalities could be explored, and although several have already been tested (see Supplemental Material Sec. V), other choices might offer even stronger bounds. Similarly, while we have tested multiple states to construct the lower bound on the p-value, we cannot rule out the possibility that better states exist for this purpose. Moreover, our statistical method could be extended to quantify finite-size effects on other squeezing or entanglement criteria [32]. We leave these extensions for future work. Nonetheless, our results already offer a solid framework for predicting and accurately assessing the outcomes of future experiments focusing on spin squeezing.